# CATNet: A geometric deep learning approach for CAT bond spread prediction in the primary market

### v1

@@ Abstract
Traditional models for pricing catastrophe (CAT) bonds struggle to capture the complex, relational data inherent in these instruments. This paper introduces CATNet, a novel framework that applies a geometric deep learning architecture, the Relational Graph Convolutional Network (R-GCN), to model the CAT bond primary market as a graph, leveraging its underlying network structure for spread prediction. Our analysis reveals that the CAT bond market exhibits the characteristics of a scale-free network, a structure dominated by a few highly connected and influential hubs. CATNet demonstrates high predictive performance, significantly outperforming a strong Random Forest benchmark. The inclusion of topological centrality measures as features provides a further, significant boost in accuracy. Interpretability analysis confirms that these network features are not mere statistical artifacts; they are quantitative proxies for long-held industry intuition regarding issuer reputation, underwriter influence, and peril concentration. This research provides evidence that network connectivity is a key determinant of price, offering a new paradigm for risk assessment and proving that graph-based models can deliver both state-of-the-art accuracy and deeper, quantifiable market insights.
@@ 1 Introduction
Traditional ML models like linear regression, decision trees, and some ensemble methods (like random forest) struggle with datasets containing categorical variables with many unique values. Each unique combination of categories (like state and peril) effectively becomes a new feature, drastically increasing the dimensionality of the dataset and making it difficult to train a robust model. Furthermore, traditional ML models struggle to represent the complex relationships between different entities involved in CAT bond contracts, such as issuers, perils, and regions. This limitation hinders their ability to accurately predict bond performance and assess risk.
@@
In this paper, we introduce CATNet, a novel framework that applies a Relational Graph Convolutional Network (R-GCN) to this domain, making several key contributions. We demonstrate that CATNet significantly outperforms the state-of-the-art Random Forest model in spread prediction using only the raw information within the bond contracts, which suggests the market is largely efficient and that the primary challenge lies in effectively representing data complexity. Our analysis also reveals that the CAT bond market has a scale-free network structure, a critical insight into its organization and potential systemic vulnerabilities. Furthermore, we provide interpretable results, showing how network centrality measures act as quantitative proxies for long-held industry intuition about issuer reputation and peril concentration. Ultimately, this work establishes a new paradigm for analyzing CAT bond primary market by prioritizing the learning of relational structure over traditional feature engineering and data manipulation.
@@ 3 Graph representation of CAT data
The power of graph formalism lies not only in its focus on relationships between points, but also in its generalization. Unlike traditional data representations that emphasize individual data points, graphs prioritize the connections between them. This focus on relationships allows for a more nuanced understanding of complex systems. Beyond providing an elegant theoretical framework, graphs offer a robust mathematical foundation for analyzing, understanding, and learning from complex systems. Before delving into representational learning on graphs, we must first establish a more formal definition of a “graph” and understand its topology.
@@ 3.2 Graph construction
Node Feature Enhancement: To enhance node representations, we integrate additional contextual features from the dataset. Specifically, we assign all remaining CAT contract information, such as the S&P rating of the bond, spread premium, expected loss, issue year, etc. (see Table 5 and 9 for the full list), to the contractID node type. This inclusion ensures that nodes encapsulate important characteristics that may influence predictions and analyses.
This comprehensive graph structure, with its rich representation of entities, relationships, and features, forms the foundation for our subsequent analysis and risk premium prediction, which will be conducted as a node-level task using a graph neural network (see Section 4). Figure 2 displays the CAT bond network for contracts issued in 2021, with node colors and edge colors indicating different types of nodes and relationships, respectively. The full graph, which contains $1902$ nodes and $8476$ edges is represented in Figure 13.
Figure 2: CAT bond network visualization for bonds issued in 2021.
@@ 3.3 CAT bond network properties
Analysis of the constructed CAT bond network offers key insight into the structure of the primary market. We begin by examining the degree distribution of the nodes to determine if the relationships within the market are random  [^2] A random graph is a graph whose nodes are connected in a random manner and as a result it has a random number of edges, refer to Appendix C.1 for more technical detail.. The degree of a node, defined as the number of edges connected to it, indicates the number of direct relationships a node has with other nodes in the graph. Figure 3 illustrates the highly skewed degree distribution of the CAT bond network, with most nodes exhibiting between 0 and 50 connections. However, the presence of several highly connected nodes suggests that the network may follow a power-law distribution, a characteristic often observed in real-world networks. We further investigate this possibility in the subsequent analysis.
@@
To study the dynamics of the CAT bond network over time, we calculate the fitness of each node, which measures the tendency of new nodes to connect to an existing node in the network  [^5] In network science, the tendency for new nodes to in a network to connect themselves to existing nodes that already have high degree (i.e., nodes that are well connected) is called preferential attachment.. Higher fitness indicates a greater likelihood that a node’s degree will increase as the network grows. The dynamics of the CAT network can be observed by analyzing sub-graphs that are expanded based on the “issue year”, starting from 1999 through 2021. For each year (i=1999,,2021)(i=1999,…,2021), we consider a sub-graph corresponding to the network of CAT contracts with an “issue year” of $i$. By continuing this process for each subsequent year, we can track the network’s evolution over the entire period.
@@
Number of Edges ($L$) 8476
Number of Nodes ($N$) 1902
@@ 4.1 Graph neural networks
where $\mathbf{h}_{u}^{(k)}$ represents the hidden state (embedding) of node $u$ at layer $k,\mathcal{N}(u)$ denotes the neighborhood of node $u,\operatorname{AGGREGATE\leavevmode\nobreak\ }^{(k)}$ is the function that aggregates messages from neighbors, and $\mathrm{UPDATE}^{(k)}$ is the function that updates the node’s embedding. Initialization is performed as $\mathbf{h}_{u}^{(0)}=\mathbf{x}_{u}$ where $\mathbf{x}_{u}$ is the initial feature vector of node $u$. This framework ensures that the learned representations are permutation invariant and effectively capture local graph structures.
@@ 4.2.1 Data transformation
4.2.2 Topological feature engineering
As demonstrated in Section 3.3, certain entities exert varying degrees of influence within the CAT bond network, suggesting that they may also influence risk premiums. To capture this influence, we generate six relevant topological features in the form of centrality measures [^10] We do not generate centrality measures for the CAT contract node, as it is not an actual entity in the network.. Betweenness Centrality measures the extent to which a node lies on paths between other nodes. Degree Centrality shows the proportion of nodes to which a node is connected. Eigenvector Centrality reflects the influence of a node on its connections. Katz Centrality is similar to Eigenvector Centrality, but accounts for the total number of walks between nodes. The clustering coefficient indicates how close a node’s neighbors are to forming a complete graph. Closeness Centrality represents the inverse of the average length of the shortest paths to all other nodes.
These features enrich the node representations and provide insights from the network structure. Figure 9 displays the top five entities in terms of their influence, as measured by each of the five centrality measures  [^11] Clustering coefficients are omitted from this plot as they are a property of a node’s neighborhood, not a unique, rankable value attributable to the entity itself.. For a more formal definition of these centrality measures, refer to Appendix C.1.
@@ 4.2.3 Training: inductive versus inductive learning paradigms on graphs
Given this context, we adopt a hybrid testing strategy. For out-of-sample (OOS) evaluation, we use a strictly transductive setup. For out-of-time (OOT) evaluation, we employ a semi-inductive variant where new edges (representing new bonds) can form between the fixed set of existing nodes. This approach allows us to probe the model’s ability to generalize to new transactions without fundamentally altering the node embedding matrix. This strategy exploits the relational power of the R-GCN architecture while acknowledging its limitations in fully inductive scenarios where entirely new entities might appear.
@@ 5 Results & Discussion
This section evaluates the performance of the R-GCN model in predicting catastrophe bond spreads against a Random Forest (RF) benchmark, an established model for tabular data (Makariou et al., 2021). Our central hypothesis is that the R-GCN’s graph-based structure can more effectively capture the complex, relational information inherent in bond contracts than a tabular model. To test this and identify key performance drivers, we conduct ablation studies by including or excluding network-derived features. We also discuss in detail the features that drive the model’s prediction.
@@ 5.1 Out-of-sample performance evaluation
We generated 10 random subsets of the data, each with an 80% training and 20% testing split. The models were trained and tested on each subset, and the average performance is reported in Table 3.
Table 3: Out of sample (OOS) prediction accuracy of R-GCN versus RF
Subset RF R2 (%) R-GCN R2 (%)
with Topo Features without Topo Features
1 66.69 74.12 63.48
2 48.55 81.52 66.77
3 47.62 72.46 59.62
4 51.48 80.26 70.93
5 44.90 74.64 68.43
6 43.42 62.87 66.96
7 47.81 79.92 60.96
8 55.79 71.95 68.57
9 48.60 75.88 62.07
10 52.5 77.69 69.00
Average 50.74 75.13 65.68
The results in Table 3 reveal a clear, two-fold advantage of the graph-based approach. We analyze these findings in two steps.
1. Impact of graph representation
First, to ensure a fair comparison based purely on the underlying data representation, we compare the RF model against the R-GCN without its supplemental topological features. In this direct comparison, the R-GCN achieves an average R2 of 65.68%, a significant improvement of over 15 percentage points compared to the RF’s average of 50.74%. This outcome strongly supports our central hypothesis: representing the bond contracts as a graph network inherently unlocks predictive power that is inaccessible to a standard tabular model.
2. Impact of topological features
Next, we demonstrate a key advantage of the deep learning graph paradigm: the ability to seamlessly engineer relevant features from the network’s structure. By incorporating six topological features (as shown in Figure 10), the R-GCN’s performance is further enhanced. The average R2 increases by nearly 10 percentage points from 65.68% to 75.13%. This highlights that critical information about market structure and participant influence can be captured “for free” as topological features providing a significant performance boost at no additional data collection cost.
@@ 5.2 Out-of-time performance evaluation
To further assess the model’s ability to generalize over time, we evaluated it using the semi-inductive framework in an out-of-time (OOT) test. The model was trained sequentially on all data up to a given year and then evaluated on the subsequent, entirely unseen year. As shown in Table 4, the R-GCN demonstrates robust temporal generalization, achieving an average R² of 72.90%. This strong performance, despite the R-GCN’s inherently transductive architecture, validates our argument from Section 4.2.3: because the CAT bond market consists of a finite set of recurring entities, the model can effectively leverage learned entity embeddings to make accurate predictions on new transactions.
The year-over-year performance exhibits some volatility, which can be attributed to varying data availability. For instance, the lower R2 scores in 2019 and 2020 correspond to years with fewer bond issuances, resulting in smaller test sets. This phenomenon, where performance evaluation can be skewed by limited data instances in a given period, has been observed in prior work (Makariou et al., 2021) and reflects the natural hard and soft cycles of the reinsurance market.
This performance variability also suggests the presence of temporal regime shifts within the market. Our own feature analysis supports this, revealing that temporal features, specifically the issuance date decomposed into sinusoidal components (see Figure 10), are highly significant predictors. While our current model captures this implicitly, a promising avenue for future research would be to employ dynamic temporal graph neural networks. Such models are explicitly designed to capture the evolving nature of graph structures and could potentially model these market regime dependencies more directly.
Table 4: Out of time (OOT) prediction accuracy of R-GCN using different subgraphs by issuance year.
Train Years Test Year R2 (%)
1999–2015 2016 79.47
1999–2016 2017 84.78
1999–2017 2018 67.94
1999–2018 2019 65.88
1999–2019 2020 56.94
1999–2020 2021 82.39
Average 72.90
@@
In summary, the results show that tabular models like Random Forest, even when given the same initial contract information, may not adequately exploit the complex web of relationships within catastrophe bond data. The graph representation itself provides a foundational performance lift, and the subsequent inclusion of network-derived features offers a second, significant improvement, leading to a state-of-the-art predictive model.

### v2

@@ Abstract
Traditional models for pricing catastrophe (CAT) bonds struggle to capture the complex, relational data inherent in these instruments. This paper introduces CATNet, a novel framework that applies a geometric deep learning architecture, the Relational Graph Convolutional Network (R-GCN), to model the CAT bond primary market as a graph, leveraging its underlying network structure for spread prediction. Our analysis reveals that the CAT bond market exhibits the characteristics of a scale-free network, a structure dominated by a few highly connected and influential hubs. CATNet demonstrates higher predictive performance, significantly outperforming strong Random Forest and XGBoost benchmarks. Interpretability analysis confirms that the network’s topological properties are not mere statistical artifacts; they are quantitative proxies for long-held industry intuition regarding issuer reputation, underwriter influence, and peril concentration. This research provides evidence that network connectivity is a key determinant of price, offering a new paradigm for risk assessment and proving that graph-based models can deliver both state-of-the-art accuracy and deeper, quantifiable market insights.
@@ 1 Introduction
Traditional ML models like linear regression, decision trees, and some ensemble methods (like random forest and XG Boost) struggle with datasets containing categorical variables with many unique values. Each unique combination of categories (like state and peril) effectively becomes a new feature, drastically increasing the dimensionality of the dataset and making it difficult to train a robust model. Furthermore, traditional ML models struggle to represent the complex relationships between different entities involved in CAT bond contracts, such as issuers, perils, and regions. This limitation hinders their ability to accurately predict bond performance and assess risk.
@@
In this paper, we introduce CATNet, a novel framework that applies a Relational Graph Convolutional Network (R-GCN) to this domain, making several key contributions. CATNet significantly outperforms state-of-the-art tabular models (Random Forest and XGBoost) in spread prediction using only raw bond contract data. This suggests that the market is largely efficient and that the primary predictive challenge lies in effectively representing data complexity.
Our analysis also reveals that the CAT bond market has a scale-free network structure, a critical insight into its organization and potential systemic vulnerabilities. Furthermore, we provide interpretable results, showing how network centrality measures act as quantitative proxies for long-held industry intuition about issuer reputation and peril concentration. Ultimately, this work establishes a new paradigm for analyzing CAT bond primary market by prioritizing the learning of relational structure over traditional feature engineering and data manipulation.
@@ 3 Graph representation of CAT data
The power of graph formalism lies not only in its focus on relationships between points, but also in its generalization. Unlike traditional data representations that emphasize individual data points, graphs prioritize the connections between them. This focus on relationships allows for a more nuanced understanding of complex systems.
Beyond providing an elegant theoretical framework, graphs offer a robust mathematical foundation for analyzing, understanding, and learning from complex systems. Before delving into representational learning on graphs, we must first establish a more formal definition of a “graph” and understand its topology.
@@ 3.2 Graph construction
Node Feature Enhancement: To enhance node representations, we integrate additional contextual features from the dataset. Specifically, we assign all remaining CAT contract information, such as the S&P rating of the bond, spread premium, expected loss, issue year, etc. (see Tables 59 for the full list), to the contractID node type. This inclusion ensures that nodes encapsulate important characteristics that may influence predictions and analyses.
The resulting graph structure, containing 1,095 nodes and 49,464 edges (Figure 13), provides the foundation for our node-level analysis. Figure 2 displays the CAT bond network for contracts issued in 2021. Although the dataset originates from 803 catastrophe bond contracts, the graph transformation creates an expansive relational environment that provides ample data for a deep learning model. This approach effectively addresses potential concerns regarding sample size; whereas a deep model would inevitably overfit on a tabular dataset of 803 rows, the R-GCN thrives on the rich topology of the graph. By shifting the learning task from isolated rows to interconnected entities, the model can extract nuanced features from the nearly 50,000 edges, ensuring high predictive performance without compromising model stability.
Figure 2: CAT bond network visualization for bonds issued in 2021. Larger nodes indicate higher degrees. Different edge colors signify unique relationships between the entities.
@@ 3.3 CAT bond network properties
Analysis of the constructed CAT bond network offers key insight into the structure of the primary market. We begin by examining the degree distribution of the nodes to determine if the relationships within the market are random  [^2] A random graph is a graph whose nodes are connected in a random manner and as a result it has a random number of edges, refer to Appendix C.1 for more technical detail.. The degree of a node, defined as the number of edges connected to it, indicates the number of direct relationships a node has with other nodes in the graph.
Figure 3 illustrates the highly skewed degree distribution of the CAT bond network, with most nodes exhibiting between 0 and 50 connections. However, the presence of several highly connected nodes suggests that the network may follow a power-law distribution, a characteristic often observed in real-world networks. We further investigate this possibility in the subsequent analysis.
@@
To study the dynamics of the CAT bond network over time, we calculate the fitness of each node, which measures the tendency of new nodes to connect to an existing node in the network  [^5] In network science, the tendency for new nodes to in a network to connect themselves to existing nodes that already have high degree (i.e., nodes that are well connected) is called preferential attachment.. Higher fitness indicates a greater likelihood that a node’s degree will increase as the network grows.
The dynamics of the CAT network can be observed by analyzing sub-graphs that are expanded based on the “issue year”, starting from 1999 through 2021. For each year (i=1999,,2021)(i=1999,…,2021), we consider a sub-graph corresponding to the network of CAT contracts with an “issue year” of $i$. By continuing this process for each subsequent year, we can track the network’s evolution over the entire period.
@@
Number of Edges ($L$) 49464
Number of Nodes ($N$) 1095
@@ 4.1 Graph neural networks
where $\mathbf{h}_{u}^{(k)}$ represents the hidden state (embedding) of node $u$ at layer $k,\mathcal{N}(u)$ denotes the neighborhood of node $u,\operatorname{AGGREGATE\penalty 10000\ }^{(k)}$ is the function that aggregates messages from neighbors, and $\mathrm{UPDATE}^{(k)}$ is the function that updates the node’s embedding. Initialization is performed as $\mathbf{h}_{u}^{(0)}=\mathbf{x}_{u}$ where $\mathbf{x}_{u}$ is the initial feature vector of node $u$. This framework ensures that the learned representations are permutation invariant and effectively capture local graph structures.
@@ 4.2.1 Data transformation
4.2.2 Quantifying structural influence via topological diagnostics
A central premise of CATNet is that by representing the catastrophe bond market as a relational graph, we eliminate the need for manual feature engineering and external macroeconomic data. In this paradigm, the model’s predictive power stems from the inherent topology of the network itself. Unlike tabular models that treat contracts as independent observations, the R-GCN architecture captures the “connectedness” of the market participants through its message-passing mechanism.
To illustrate the quantitative significance of this structural information, we use six centrality measures as topological diagnostics (see Figure 9). It is important to distinguish these from traditional feature engineering; these measures are not external inputs but are mathematical byproducts of the graph’s existing node-edge relationships. By explicitly including these metrics in our node importance analysis (see Section 5.3.3, Figure 10), we can “unmask” the latent relational patterns that the R-GCN exploits.
For diagnostic purposes, we initially employed these features in a unidirectional graph setting, where they provided a significant performance boost by compensating for limited relational flow. In our more advanced bidirectional implementation, we observed that these explicit features became redundant, as the model’s enhanced message-passing layers learned the equivalent structural information autonomously. We include them in our reported feature importance rankings to provide transparency into the “black box” of the GNN, demonstrating that market influence—proxied by metrics like Betweenness and Closeness Centrality—is a statistically significant determinant of CAT bond spreads. For a more formal definition of these centrality measures, refer to Appendix C.1
@@ 4.2.3 Training: inductive versus inductive learning paradigms on graphs
We employ a hybrid testing strategy for out-of-sample (OOS) and out-of-time (OOT) evaluations based on a transductive setup with masked test features. To ensure the model learns from the graph’s topology rather than relying on feature leakage, the feature vectors of test nodes are zeroed out during training while their structural positions are preserved. At evaluation, the true features are reintroduced, and the model predicts spreads using frozen weights. This methodology allows us to evaluate the model’s generalization capabilities without altering the underlying node embedding matrix. By using this approach, we leverage the relational strengths of the R-GCN architecture while transparently addressing its limitations in fully inductive environments where entirely new entities might appear.
@@ 5 Results & Discussion
This section evaluates the performance of the R-GCN model in predicting catastrophe bond spreads against two robust tabular benchmarks: Random Forest (RF) (Makariou et al., 2021) and XGBoost (Chen et al.,, 2024). Our central hypothesis is that the R-GCN’s graph-based structure can more effectively capture the complex, relational information inherent in bond contracts than a tabular model. To test this and identify key performance drivers, we conduct ablation studies by including or excluding network-derived features. We also discuss in detail the features that drive the model’s prediction.
@@ 5.1 Out-of-sample performance evaluation
We generated 10 random subsets of the data, each with an 80% training and 20% testing split. The models were trained and tested on each subset, and the average performance is reported in Table 3. The R-GCN model consistently outperformed established tabular benchmarks, achieving the highest average $R^{2}$ of 79.05%, compared to 77.63% for XGBoost and 73.09% for Random Forest. Notably, the R-GCN recorded the lowest average error (MSE: 0.2109), representing a significant 22.0% reduction in error relative to the Random Forest baseline and a 7.7% reduction compared to XGBoost.
The R-GCN’s superior performance is attributed to its ability to leverage the market’s underlying relational structure—specifically the connections between bonds sharing common market participants or peril types. This structural information serves as an implicit regularizer, allowing the model to generalize more effectively across the heterogeneous CAT bond universe than models limited to independent feature mappings. While tree-based models like XGBoost may remain competitive in instances where relational connectivity is sparse, the R-GCN provides a more robust framework for capturing the complex interdependencies of the primary market.
Table 3: OOS model performance across folds
Fold Random Forest XGBoost R-GCN
MSE $\mathbf{R^{2}}$ (%) MSE $\mathbf{R^{2}}$ (%) MSE $\mathbf{R^{2}}$ (%)
1 0.4345 61.21 0.3588 67.97 0.2754 75.42
2 0.1680 79.74 0.1573 81.02 0.1088 86.88
3 0.3051 73.92 0.2506 78.58 0.2849 75.64
4 0.2218 73.69 0.2024 75.99 0.2472 70.67
5 0.2199 64.72 0.1422 77.19 0.1231 80.26
6 0.3610 74.09 0.2737 80.36 0.2419 82.64
7 0.2262 80.72 0.2605 77.81 0.1958 83.31
8 0.3275 69.91 0.3013 72.32 0.2743 74.80
9 0.1842 75.79 0.1101 85.52 0.1671 78.03
10 0.2549 77.07 0.2275 79.54 0.1907 82.85
Average 0.2703 73.09 0.2284 77.63 0.2109 79.05
@@ 5.2 Out-of-time performance evaluation
Table 4 reports performance under the out-of-time protocol, where models are trained on historical data and evaluated on future test years. This walk-forward design mirrors realistic deployment conditions requiring models to generalize across shifting market regimes.
The R-GCN achieves a superior average $R^{2}$ of 76.79%, surpassing XGBoost (68.88%) and Random Forest (66.28%) by 7.91 and 10.51 percentage points (pp), respectively. Its average MSE of 0.1511 represents a 24.1% reduction relative to XGBoost. Notably, the R-GCN’s improvement margin over tabular baselines is substantially larger in the OOT setting than in the OOS context (7.91pp vs. 1.42pp), suggesting that relational graph structures are particularly robust for temporal extrapolation.
The R-GCN outperforms benchmarks in five of the six temporal splits, reaching its peak $R^{2}$ in 2021 (89.03%). Its most significant advantage occurs in 2018, where its $R^{2}$ of 80.04% exceeds XGBoost’s 53.48% by a margin of 26.56pp. Although XGBoost leads the 2019 cohort (72.97% vs. 63.14%), the R-GCN retains its superior performance during the 2020 COVID-19 market disruption ($R^{2}$: 51.77%).
This enhanced generalization is attributed to the message-passing mechanism, which propagates pricing signals through stable entity-level relationships—including cedents, underwriters, and perils—that persist across market cycles. While individual bond features fluctuate across regimes, the relational topology provides a time-invariant learning signal that tabular models, operating on isolated feature vectors, cannot exploit.
Table 4: OOT model performance across walk-forward splits
Train / Val Split Test Year Random Forest XGBoost R-GCN
MSE $\mathbf{R^{2}}$ (%) MSE $\mathbf{R^{2}}$ (%) MSE $\mathbf{R^{2}}$ (%)
1999–2014 — Val 2015 2016 0.1722 74.92 0.1205 82.45 0.0820 88.05
1999–2015 — Val 2016 2017 0.1578 73.39 0.1149 80.63 0.0668 88.73
1999–2016 — Val 2017 2018 0.2335 62.01 0.2860 53.48 0.1227 80.04
1999–2017 — Val 2018 2019 0.2295 69.13 0.2010 72.97 0.2741 63.14
1999–2018 — Val 2019 2020 0.3252 45.01 0.3237 45.27 0.2853 51.77
1999–2019 — Val 2020 2021 0.1851 73.23 0.1489 78.46 0.0758 89.03
Average 0.2172 66.28 0.1992 68.88 0.1511 76.79
@@
We validated the assumption that macroeconomic information is already embedded in catastrophe bond prices at issuance by introducing two market indicators: the BBB corporate credit spread and the Guy Carpenter Rate-on-Line (see Appendix B.2, Tables 1112). As summarized in our findings, these macroeconomic covariates produced asymmetric effects across model architectures. In the OOS setting, Random Forest and XGBoost improved slightly (+2.57% and +0.47%, respectively), while the R-GCN declined by 2%. The out-of-time (OOT) setting revealed a sharper divide: while Random Forest maintained a 2.39% gain, XGBoost deteriorated significantly (-4.34%) and the R-GCN declined further (-2.51%).
These divergent responses reflect architectural differences in covariate processing. Random Forest’s bagging and feature sub-sampling naturally regularize against uninformative inputs. Conversely, XGBoost’s sequential boosting aggressively fits residual patterns, yielding modest OOS gains but substantial OOT degradation as market regimes shift—a byproduct of the non-stationary nature of credit spread dynamics.
The R-GCN’s sensitivity arises from its message-passing mechanism. Because bonds issued within the same year share temporal edges, macroeconomic features—which are identical for the entire cohort—are reinforced through neighborhood aggregation. This amplification causes the model to overweight uniform temporal signals at the expense of primary pricing drivers like expected loss or maturity. Unlike tabular models that score observations independently, the graph structure transforms these covariates into a network-wide shift in learned representations.
In summary, the results show that tabular models, even when given the same initial contract information, may not adequately exploit the complex web of relationships within catastrophe bond data. The graph representation itself provides a foundational performance lift, and the subsequent inclusion of network-derived features offers a second, significant improvement, leading to a state-of-the-art predictive model.